# GenEOL: Harnessing the Generative Power of LLMs for Training-Free Sentence Embeddings

### v1

@@ Abstract
Training-free embedding methods directly leverage pretrained large language models (LLMs) to embed text, bypassing the costly and complex procedure of contrastive learning. Previous training-free embedding methods have mainly focused on optimizing embedding prompts and have overlooked the benefits of utilizing the generative abilities of LLMs. We propose a novel method, GenEOL, which uses LLMs to generate diverse transformations of a sentence that preserve its meaning, and aggregates the resulting embeddings of these transformations to enhance the overall sentence embedding. GenEOL significantly outperforms the existing training-free embedding methods by an average of 2.85 points across several LLMs on the sentence semantic text similarity (STS) benchmark. Our analysis shows that GenEOL stabilizes representation quality across LLM layers and is robust to perturbations of embedding prompts. GenEOL also achieves notable gains on multiple clustering, reranking and pair-classification tasks from the MTEB benchmark.
@@
Raghuveer Thirukovalluru1, Bhuwan Dhingra1, 1Duke University, rt195@duke.edu
@@ 1 Introduction
We specifically focus on the semantic text similarity (STS) benchmark for sentence embeddings Conneau and Kiela (2018). Our key contributions are as follows: 1. We show that diverse meaning retaining transformations are helpful in improving training free embeddings. 2. The proposed method, GenEOL, at a higher capacity ($m$=32), significantly outperforms the previous best training-free method by (3.88, 1.83, 2.83) points on average with (Mistral0.1-7B, Llama2-7B, Llama3-8B) respectively. GenEOL with even a small number of transformed sentences ($m$=2) surpasses all baselines. 3. GenEOL stabilizes representational quality across different layers of LLMs. GenEOL is robust to changes in embedding prompts. 4. GenEOL also performs better than baselines across multiple diverse MTEB tasks Muennighoff et al. (2022).
@@ 2 Background and Related Work
Generating data for CL Training: E5 Wang et al. (2023) used ChatGPT to generate a huge CL corpus of related positives and unrelated negatives across multiple tasks. Gecko Lee et al. (2024) used an LLM to generate queries and relabel positives, negatives from existing data corpus. LLM2Vec BehnamGhader et al. (2024) used representations generated from different dropout masks as positives with a positive only CL loss. For sentences specifically, SynCSE Zhang et al. (2023a) defined specific transformations with ChatGPT to develop contrastive positives, negatives for sentence embedding training. Inspired by compositional transformations in CL data for computer vision (CV), SumCSE Thirukovalluru et al. (2024b) further improved these transformations by using ‘summary’ operations (akin to cropping in CV). Our method, GenEOL, takes inspiration from SynCSE, SumCSE in generating transformations (with same core meaning) to improve embedding quality.
@@
Training Free Methods: Echo Springer et al. (2024) showed that repeating the original sentence and using the hiddden representations of the later as the embedding of the sentence can improve performance. PromptEOL Jiang et al. (2023) showed that generative LLMs when used with the prompt (This sentence: "[TEXT]" means in one word:") and using the hidden representation of the last token i.e. " is very effective at embedding the text. Further improving this prompt Zhang et al. (2024) proposed two advanced prompts - 1. Pretended Chain of Thought prompt (PCoTEOL) - (After thinking step by step , this sentence : “[X]” means in one word:“) and 2. Knowledge Enhancement prompt (KEEOL) - (The essence of a sentence is often captured by its main subjects and actions….“[X]” means in one word:“). MetaEOL Lei et al. (2024) utilized eight diverse prompts from four categories—Text Classification, Sentiment Analysis, Paraphrase Identification, and Information Extraction—to aggregate multiple perspectives into a single final embedding. These methods also discuss better ways to extract text embeddings by using LLM penultimate layers. However, they do not fully leverage the generative power of LLMs for embedding sentences.
@@ 3.1 Motivation
Observation 1: Instruction tuned LLMs are worse than Pretrained LLMs at Embeddings We first analyze in Fig. 2(a), the performance (spearman rank correlation) of instruction-tuned LLMs versus pretrained models on the STSB validation set using the KEEOL prompt Zhang et al. (2024) for sentence encoding. Results indicate that pretrained models consistently outperform instruction-tuned models in embeddings. Additionally, when applying an instruction version of the knowledge EOL prompt (with [INST] tokens) to instruction-tuned models, performance decreased further. This may explain why recent studies Zhang et al. (2024) and Lei et al. (2024) benchmark non-CL training-free methods primarily using pretrained models. On the other hand, instruction tuned models are generally vastly superior at generating texts suited to various prompts. This raises the question – how can we utilize the generative ability of an instruction-tuned model to enhance the embeddings obtained from the pretrained model?
@@
To reduce the spread around the diagonal, let’s consider the following exercise. Let $(x_{i1},x_{i2})$ be the i-th datapoint. Let $\tau_{i}$ be the true score and $\rho_{i}$ be the predicted similarity score of this datapoint. Considering $\rho_{i}$ to be a random variable, the coarse alignment trend from Fig 2(b) gives a sense of an error being present in $\rho_{i}$. Specifically, $\rho_{i}=\tau_{i}+\epsilon_{i}$, where $\epsilon_{i}$ is the error term with a mean $\mu_{i}$ and variance $\sigma^{2}_{i}$. From a statistics standpoint, the primary task now is to the reduce this error to improve performance. $\mu_{i}$ here is the inherent bias and would require further tuning the model to reduce it. To reduce the variance $\sigma^{2}_{i}$, a simple trick now would be to average $k$ independent estimates of $\rho_{i}$ i.e. $\{\rho^{1}_{i},..,\rho^{k}_{i}\}$. It is known that the variance of the mean of $k$ independent random variables decreases by a factor of $k$ and hence reduces the error $\epsilon_{i}$.
@@ 3.2.1 Diverse Transformations
Method Generator $\mathbf{m}$ STS12 STS13 STS14 STS15 STS16 STSB Sick-R Avg.
@@ 3.2.2 Further Increasing Diversity (Optional)
SumCSE Thirukovalluru et al. (2024b) showed that compositional summary transformations (a second summary transformation over the first transformation) is very effective at creating transformed sentences far from the original sentence. Compositional transformations retain the original meaning of the sentence while more significantly altering its lexical form compared to a single transformation. As shown in Table 1, a compositional summary transformation (§A.4) (emphasizes for meaning preservation in contrast to SumCSE prompt) shows benefits. Thus, we optionally incorporate the compositional summary transformation in our approach. The proposed framework, GenEOL, also allows for use of other diverse transformations or sampling multiple diverse sentences from these proposed transformations to improve sentence embeddings.
@@ 3.2.3 Final Embedding
The transformed sentences are individually embedded using the embedder, LLM $\mathcal{L}_{PT}$ following previous work Zhang et al. (2024). This mean embedding becomes the final embedding of the sentence.
@@ 4 Experiments
We evaluate GenEOL and other baselines on 7 semantic text similarity tasks - STS12, STS13, STS14, STS16, STS16, STSB, SICK-R Gao et al. (2021). Spearman rank correlation (cosine similarity) is the main metric Muennighoff et al. (2022). Training sets of the STS tasks are not used. We additionally asses GenEOL on 10 MTEB tasks across 4 categories (Classification, Clustering, Reranking and Pair Classification) using the appropriate metrics described in Muennighoff et al. (2022).
@@ 4.1 Implementation
Embedder in GenEOL, by default, uses the embedding prompt from KEEOL Zhang et al. (2024) i.e. "The essence of a sentence is often captured by its main subjects and actions, while descriptive terms provide additional but less central details. With this in mind , this sentence : “[X]” means in one word:“". We evaluate the performance of GenEOL at different values of $m$. When $m>4$, multiple transformed sentences are sampled from the generator from each transformation.
4.2 Main Table
@@
4.3 Ablation 1: Effect of increasing the number of transformed sentences, $m$
The number of transformed sentences, denoted as $m$, is a key parameter that significantly influences the performance of GenEOL. To provide a clearer understanding of GenEOL’s effectiveness, we present results with varying $m$ values for both the Mistral0.1-7B and Llama3-8B embedding models in Fig. 3. For $m=2$, we randomly sample from transformations $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$. For $m\geq 4$, we maintain and equal diversity across all transformations - ($\mathcal{T}_{1}...\mathcal{T}_{4}$). Even with just two generations, GenEOL ($m=2$) beats all other baselines.
Increasing $m$ results in a significant increase in performance from $m=0$ to $m=32$ of over 5 points. The improvements are lower at higher $m$. Performance starts to stagnate after $m=16$. This is because similar transformed sentences/repetitions are sampled from the generator at higher $m$ for a fixed number of transformations (4 in our case). Hence, more diverse transformations might be required at higher $m$ values.
@@ 4.4 Ablation 2: Dissecting GenEOL
Figure 3: Average STS performance with $m$. GenEOL beats MetaEOL with only 2 transformations ($m$=2) for both embedders. Performance starts to stagnate around 16 generations. Results averaged over 3 seeds.
@@ 4.5 Ablation 3: Effect of Layer Number
Prior work on training free embeddings has shown that the last hidden layer might not be the most appropriate layer for text embeddings Lei et al. (2024); Zhang et al. (2024); Li and Li (2024). We hence evaluate the performance of GenEOL across different hidden layers. We perform the same analysis for KEEOL. Penultimate layer does best for Llama3-8B and second best for Mistral0.1-7B. Performance drops beyond the penultimate layer.
@@
4.6 Compute Allocation in GenEOL
@@
Method Classification Clustering Reranking Pair Classification
AC B7 EC MX TN AU SD SO TS SD
KEEOL 72.72 78.70 48.46 24.54 30.48 53.95 73.44 39.12 71.03 43.70
TSEOL 74.91 82.89 50.70 30.88 50.87 62.52 85.26 47.78 73.75 69.18
GenTSEOL($m=8$) 71.59 82.73 47.54 31.14 51.78 63.58 85.44 48.45 76.82 77.58
4.7 Sensitivity to Embedder Prompts
4.8 Sensitivity to Generator Prompts
Qualitative Analysis: Some examples for each of the four runs from Table 7 are shown in §A.6. Poor quality generations that missed out essential details/hallucinated new content are highlighted. As can be seen, Llama-I-3.1 with few shot prompts and Mistral0.1-I-7B with zero shot prompt produced poor transformations.
@@ 4.8 Sensitivity to Generator Prompts
4.9 More MTEB Datasets
We use the same transformations as described in §3.2 to generate transformed sentences for the above tasks. We use ChatGPT as the generator to asses the performance, minimizing any poor generations. These transformed generations are then used with TSEOL prompts to get numbers for GenTSEOL. GenTSEOL results in notable gains across tasks except for classification.
Classification tasks often rely on specific aspects of a sentence (E.g. ‘Emotion’ in EC and ‘Intent’ in B7). These aspects are very specific to the original sentence. Transformations in GenTSEOL often change sentence structure, active speech to passive, etc. and might not retain such specific aspects like emotion. Hence, GenTSEOL performs worse on classification tasks. The proposed GenTSEOL framework allows for other transformations than ones prescribed in §3.2. We believe such task specific transformations can further improve GenTSEOL. We leave this exploration for future work.

### v2

@@ Abstract
Training-free embedding methods directly leverage pretrained large language models (LLMs) to embed text, bypassing the costly and complex procedure of contrastive learning. Previous training-free embedding methods have mainly focused on optimizing embedding prompts and have overlooked the benefits of utilizing the generative abilities of LLMs. We propose a novel method, GenEOL, which uses LLMs to generate diverse transformations of a sentence that preserve its meaning, and aggregates the resulting embeddings of these transformations to enhance the overall sentence embedding. GenEOL significantly outperforms the existing training-free embedding methods by an average of 2.85 points across several LLMs on the sentence semantic text similarity (STS) benchmark. GenEOL also achieves notable gains in clustering, reranking, and pair-classification tasks from the MTEB benchmark. Additionally, GenEOL stabilizes representation quality across LLM layers and remains robust to perturbations of embedding prompts.
@@
Raghuveer Thirukovalluru1, Bhuwan Dhingra1, 1Duke University, raghuveer.thirukovalluru@duke.edu
@@ 1 Introduction
Our key contributions are as follows: 1. We show that diverse meaning retaining transformations are helpful in improving training free embeddings. 2. The proposed method, GenEOL, at a higher capacity ($m$=32), significantly outperforms the previous best training-free method by (3.88, 1.83, 2.83) points on average with (Mistral0.1-7B, Llama2-7B, Llama3-8B) resp. on the STS benchmark Conneau and Kiela (2018). GenEOL with as few as two transformed sentences ($m$=2) surpasses all baselines. 3. GenEOL outperforms other training-free methods on diverse MTEB tasks, even surpassing strong unsupervised methods (that include contrastive training) like LLM2Vec BehnamGhader et al. (2024). 4. GenEOL stabilizes representational quality across different LLM layers. GenEOL is also robust to perturbations in embedding prompts.
@@ 2 Background and Related Work
Generating data for CL Training: E5 Wang et al. (2023) used ChatGPT to generate a huge CL corpus of related positives and unrelated negatives across multiple tasks. Gecko Lee et al. (2024) used an LLM to generate queries and relabel positives, negatives from existing data corpus. LLM2Vec BehnamGhader et al. (2024) used representations generated from different dropout masks as positives with a positive only CL loss. For sentences specifically, SynCSE Zhang et al. (2023a) defined specific transformations with ChatGPT to develop contrastive positives, negatives for sentence embedding training. Inspired by compositional transformations in CL data for computer vision (CV), SumCSE Thirukovalluru et al. (2024b) further improved these transformations by using ‘summary’ operations (akin to cropping in CV). Our method, GenEOL, takes inspiration from SynCSE, SumCSE in generating meaning retaining transformations.
@@
Training Free Methods: Echo Springer et al. (2024) showed that repeating the original sentence and using the hiddden representations of the later as the embedding of the sentence can improve performance. PromptEOL Jiang et al. (2023) showed that generative LLMs when used with the prompt (This sentence: "[TEXT]" means in one word:") and using the hidden representation of the last token i.e. :" is very effective at embedding the text. Further improving this prompt Zhang et al. (2024) proposed two advanced prompts - 1. Pretended Chain of Thought prompt (PCoTEOL) - (After thinking step by step , this sentence : "[X]" means in one word:") and 2. Knowledge Enhancement prompt (KEEOL) - (The essence of a sentence is often captured by its main subjects and actions…."[X]" means in one word:"). MetaEOL Lei et al. (2024) utilized eight diverse prompts from four categories—Text Classification, Sentiment Analysis, Paraphrase Identification, and Information Extraction—to aggregate multiple perspectives into a single final embedding. These methods also discuss better ways to extract text embeddings by using LLM penultimate layers. However, they do not fully leverage the generative power of LLMs for embedding sentences.
@@ 3.1 Motivation
Observation 1: Instruction tuned LLMs are worse than Pretrained LLMs at Embeddings We first analyze in Fig. 2(a), the performance (spearman rank correlation) of instruction-tuned LLMs versus pretrained models on the STSB validation set using the KEEOL prompt Zhang et al. (2024) for sentence encoding. Results indicate that pretrained models consistently outperform instruction-tuned models in embeddings. Additionally, when applying an instruction version of the knowledge EOL prompt (with [INST] tokens) to instruction-tuned models, performance decreased further. This may explain why recent studies Zhang et al. (2024) and Lei et al. (2024) benchmark non-CL training-free methods primarily using pretrained models. On the other hand, instruction tuned models are generally vastly superior at generating texts suited to various prompts. This motivates an investigation into the potential of leveraging the generative capabilities of an instruction-tuned model $\mathcal{L}_{IT}$ to enhance embeddings derived from a pretrained model $\mathcal{L}_{PT}$.
@@
To reduce the spread around the diagonal, we consider the following exercise. Let $(x_{i1},x_{i2})$ be the i-th datapoint. Let $\tau_{i}$ be the true score and $\rho_{i}$ be the predicted similarity score of this datapoint. Considering $\rho_{i}$ to be a random variable, the coarse alignment trend from Fig 2(b) gives a sense of an error being present in $\rho_{i}$. Specifically, $\rho_{i}=\tau_{i}+\epsilon_{i}$, where $\epsilon_{i}$ is the error term with a mean $\mu_{i}$ and variance $\sigma^{2}_{i}$. From a statistics standpoint, the primary task now is to the reduce this error to improve performance. $\mu_{i}$ here is the inherent bias and would require further tuning the model to reduce it. To reduce the variance $\sigma^{2}_{i}$, a simple trick now would be to average $k$ independent estimates of $\rho_{i}$ i.e. $\{\rho^{1}_{i},..,\rho^{k}_{i}\}$. It is known that the variance of the mean of $k$ independent random variables decreases by a factor of $k$ and hence reduces the error $\epsilon_{i}$.
@@
Method Generator $\boldsymbol{m}$ STS12 STS13 STS14 STS15 STS16 STSB Sick-R Avg.
@@ 3.2.2 Further Increasing Diversity (Optional)
SumCSE Thirukovalluru et al. (2024b) showed that compositional summary transformations (a second summary transformation over the first transformation) is very effective at creating transformed sentences far from the original sentence. Compositional transformations retain the original meaning of the sentence while more significantly altering its lexical form compared to a single transformation. As shown in Table 1, a compositional summary transformation (§A.4) (emphasizes for meaning preservation in contrast to SumCSE prompt) shows benefits. Thus, we optionally incorporate the compositional summary transformation.
@@ 3.2.3 Final Embedding
The proposed framework, GenEOL, allows for use of other diverse transformations or sampling multiple diverse sentences ($m$ number of them) from the proposed transformations to improve sentence embeddings. The transformed sentences ($m$) and the original sentence ($1$) are embedded using the embedder, LLM $\mathcal{L}_{PT}$ using the KEEOL prompt Zhang et al. (2024). As shown in Fig. 1, the final sentence embedding is the mean of the ($m+1$) embeddings.
@@ 4 Experiments
We evaluate GenEOL and other baselines on the STS benchmark Conneau and Kiela (2018). Spearman rank correlation (cosine similarity) is the main metric Muennighoff et al. (2022). Training sets of the STS tasks are not used. We additionally asses GenEOL on 10 MTEB tasks across 4 categories (Classification, Clustering, Reranking and Pair Classification) using the appropriate metrics described in Muennighoff et al. (2022).
@@ 4.1 Implementation
Embedder in GenEOL, by default, uses the embedding prompt from KEEOL Zhang et al. (2024) i.e. The essence of a sentence is often captured by its main subjects and actions, while descriptive terms provide additional but less central details. With this in mind , this sentence : "[X]" means in one word:". We evaluate the performance of GenEOL at different values of $m$. When $m>4$, multiple transformed sentences are sampled from the generator from each transformation.
4.2 Results STS
@@
Method Embedder Classification Clustering Reranking Pair Classification Avg.
AC B7 EC MX TN AU SD SO TS SD
Unsupervised
SimCSE Bert-Large 67.09 73.55 42.22 21.97 23.21 51.57 66.33 39.35 60.21 69.41 51.49
LLM2Vec Mistral0.1-7B 76.94 86.16 48.88 26.93 30.26 58.60 77.81 49.80 68.76 91.30 61.54
Training-Free
PromptEOL Mistral0.1-I-7B 71.91 76.70 45.75 26.52 36.27 55.85 78.54 43.00 67.25 32.21 53.40
PCoTEOL 71.88 75.89 44.10 25.58 34.17 57.30 78.27 45.10 67.82 40.24 54.03
KEEOL 72.72 78.70 48.46 24.54 30.48 53.95 73.44 39.12 71.03 43.70 53.61
MetaEOL 74.85 83.26 52.75 28.48 41.16 57.21 80.68 43.27 68.76 72.49 60.29
TSEOL 74.91 82.89 50.70 30.88 50.87 62.52 85.26 47.78 73.75 69.18 62.87
GenTSEOL ($m=8$) 71.59 82.73 47.54 31.14 51.78 63.58 85.44 48.45 76.82 77.58 63.67
4.3 Results MTEB
We use the same transformations as described in §3.2 to generate transformed sentences for the above tasks. We use ChatGPT as the generator to asses the performance, minimizing any poor generations. These transformed generations are then used with TSEOL prompts to get numbers for GenTSEOL. As shown in Table 3, GenTSEOL outperforms all other training-free methods on the MTEB tasks. It also beats strong unsupervised methods like LLM2Vec BehnamGhader et al. (2024) (which include contrastive training).
GenTSEOL results in notable gains across diverse tasks except for classification. Classification tasks often rely on specific aspects of a sentence (E.g. ‘Emotion’ in EC and ‘Intent’ in B7). These aspects are very specific to the original sentence. Transformations in GenTSEOL often change sentence structure, active speech to passive, etc. and might not retain such specific aspects like emotion. Hence, GenTSEOL performs worse on classification tasks. The proposed GenTSEOL framework allows for other transformations than ones prescribed in §3.2. We believe such task specific transformations can further improve GenTSEOL. We leave this exploration for future work.
4.4 Ablation 1: Effect of increasing the number of transformed sentences, $\boldsymbol{m}$
The number of transformed sentences, denoted as $m$, is a key parameter that significantly influences the performance of GenEOL. To provide a clearer understanding of GenEOL’s effectiveness, we present results with varying $m$ values for both the Mistral0.1-7B and Llama3-8B embedding models in Fig. 3. For $m=2$, we randomly sample from transformations $\mathcal{T}_{1}$ and $\mathcal{T}_{2}$. For $m\geq 4$, we maintain and equal diversity across all transformations - ($\mathcal{T}_{1}...\mathcal{T}_{4}$). Even with just two generations, GenEOL ($m=2$) beats all other baselines on STS.
Increasing $m$ results in a significant increase in performance from $m=0$ to $m=32$ of over 5 points. The improvements are lower at higher $m$. Performance starts to stagnate after $m=16$. This is because similar transformed sentences/repetitions are sampled from the generator at higher $m$ for a fixed number of transformations (4 in our case). Hence, more diverse transformations might be required at higher $\boldsymbol{m}$ values.
@@ 4.5 Ablation 2: Dissecting GenEOL
Figure 3: Average STS performance with $m$. GenEOL beats MetaEOL with only 2 transformations ($m$=2) for both embedders. Performance starts to stagnate around 16 generations. GenEOL results averaged over 3 seeds.
@@ 4.6 Ablation 3: Effect of Layer Number
Prior work on training free embeddings has shown that the last hidden layer might not be the most appropriate layer for text embeddings Lei et al. (2024); Zhang et al. (2024); Li and Li (2024). We hence evaluate the performance of GenEOL across different hidden layers. We perform the same analysis for KEEOL. As shown in Fig. 4, penultimate layer does best for Llama3-8B and second best for Mistral0.1-7B. Performance drops beyond the penultimate layer.
@@
4.7 Ablation 4: Compute Allocation in GenEOL
@@
4.8 Sensitivity to Embedder Prompts
4.9 Sensitivity to Generator Prompts
@@
Qualitative Analysis: Some examples for each of the four new runs from Table 7 are shown in §A.6. Poor quality generations that missed out essential details/hallucinated new content are highlighted. As can be seen, Llama-I-3.1 with few shot prompts and Mistral0.1-I-7B with zero shot prompt produced poor transformations.
@@
Method Embedder STS Avg.
MetaEOL Mistral0.1-I-7B 75.85
GenEOL ($m=8$) 79.11
GenMetaEOL ($m=8$) 81.34
Table 8: Integrating GenEOL generator transformations with MetaEOL embedder prompts yields significant gains, highlighting their complementary strengths.
4.10 Discussion: Difference from MetaEOL and other training-free methods
GenEOL is a simple method that generates meaning-preserving transformations and aggregates them to improve training-free embeddings. While it may seem similar to other methods using multiple prompts, key differences exist.
For instance, MetaEOL relies solely on prompting, while GenEOL generates transformations compatible with any prompt. MetaEOL does not use any revised versions of input sentences. It uses eight embedder prompts to get multiple embeddings capturing different aspects of a sentence (Eg. One prompt captures sentiment in it; One captures entities in the sentence; Another captures opinion; One assesses the type of the sentence, i.e., if it belongs to education/business/environment classes, etc.) GenEOL uses a single embedder prompt. It uses multiple generator prompts to produce semantically equivalent transformations. To validate their synergy, we combine GenEOL transformations with MetaEOL prompts, introducing GenMetaEOL. Table 8 shows that this combination yields significant performance gains justifying the complementary nature of the methods.
QA-SimCSE Tu et al. (2024) is another method that generates intermediate text before generating embeddings for the given input. However, the work deals with a completely different task of conditional sentence similarity (An additional condition is used to assess similarity between two sentences) and it is not training-free. GenEOL produces $m$ meaning retaining transformations (as opposed to one conditional answer in QA-SimCSE) that are all used to embed the sentence.
4.11 Discussion: Addressing the extra generation cost in GenEOL
GenEOL achieves superior performance over all training-free methods on the STS and MTEB benchmarks, demonstrating its effectiveness. However, this performance gain comes with increased computational cost. Table 9 provides a detailed comparison of the per example runtime for embedders and generators across methods, highlighting the higher overhead associated with GenEOL.
GenEOL’s computational cost is higher due to generation of multiple transformations. However, this cost is incurred solely during inference, involving only forward passes, which are more readily parallelizable (vs training-intensive methods requiring backward passes) Anil et al. (2018); McCandlish et al. (2018); Thirukovalluru et al. (2024a).
Moreover, techniques such as speculative decoding Li et al. (2024) and inference engines such as vllm Kwon et al. (2023) have made great strides in reducing the cost and latency of sampling from LLMs. Additionally, one might not need all 32 transformations to do a good job. One could adaptively scale compute for individual examples similar to MRL Kusupati et al. (2022). All of these techniques can make GenEOL more efficient.
Method Embedder Generator
PromptEOL 0.05 0.00
PCoTEOL 0.05 0.00
MetaEOL 0.24 0.00
GenEOL ($m=8$) 0.18 1.60
GenEOL ($m=2$) 0.08 0.40
Table 9: Average per example runtime (in seconds) comparison on the STS. GenEOL incurs higher cost.